OpenStudy (aaronandyson):

The marked price of an article is $4500 and the rate of tax is 8%.A shopkeeper buys the article at some discount and sells it to a customer at its marked price.If the VAT paid by the shopkeeper to the government is $72.Find: The price paid by the shopkeeper to buy the article inclusive of tax,and; The % of discount he received.

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

I'm thinking, please wait...

OpenStudy (michele_laino):

If the marked price is 4500, then is the VAT paid to governement equal to: \(4500 \cdot 8/100\)?

OpenStudy (michele_laino):

I think that the price \(x\) paid by the shopkeeper is given by the subsequent equation: \[x + 72 + \frac{8}{{100}} \cdot \left( {x + 72} \right) = 4500\]

OpenStudy (aaronandyson):

VAT to the government is given as $72

OpenStudy (michele_laino):

and \(8 \% \) is not a VAT, right?

OpenStudy (aaronandyson):

8 is rate of tax.

OpenStudy (aaronandyson):

8%*

OpenStudy (michele_laino):

I think that the right equation is the equation above, namely: \[\left( {x + 72} \right) + \frac{8}{{100}} \cdot \left( {x + 72} \right) = 4500\]

OpenStudy (aaronandyson):

What does x represent?

OpenStudy (michele_laino):

the shopkeeper buys the article to a price \(x\) and then he pays the VAT, so he spends \(x+72\) in total, right?

OpenStudy (michele_laino):

am I right?

OpenStudy (aaronandyson):

His price inclusive of tax is 4500*108/100 that's 4860

OpenStudy (michele_laino):

question: have I add the VAT to the marked price, before taxes or after taxes?

OpenStudy (aaronandyson):

VAT = tax charged - tax paid

OpenStudy (aaronandyson):

OR Vat = tax on profit

OpenStudy (michele_laino):

do you have the options?

OpenStudy (aaronandyson):

No options.

OpenStudy (michele_laino):

please follow my reasoning, and tell me if I am wrong

OpenStudy (michele_laino):

the shopkeeper buys the product to a price x, then he has to pay tax which is 8%, so he pays in total: x+ 8% x, am I right?

OpenStudy (aaronandyson):

The priice he buyed the article is given as $4500(exclusive of tax)

OpenStudy (michele_laino):

sorry he buys that article at a discount, so I think he pays a price less than 4500, right?

OpenStudy (aaronandyson):

Yes,and we have to calculate the discount

OpenStudy (michele_laino):

so he pays a price \(x\) and with tax, such price is \(x +0.08 \cdot x=1.08\;x\)

OpenStudy (michele_laino):

Now i add \(72\) so i get this equation: \((1.08 \cdot x)+72=4860\)

OpenStudy (aaronandyson):

4500 adding the tax 360 he pays 4860

OpenStudy (michele_laino):

question: is the expression \(1.08 \;x +72\) the price paid by the shopkeeper?

OpenStudy (aaronandyson):

no

OpenStudy (michele_laino):

I think that the shopkeeper pays a price of \(x+72\), so we can write the subsequent equation: \(x+72=4860\)

OpenStudy (michele_laino):

and the \(\%\) of the discount is: \((72/x) \cdot 100\)

OpenStudy (phi):

I don't fully understand the question, but using Vat = tax on profit we can say 0.08* profit = 72 profit = 72/0.08= 900 If he sells at 4500, with a profit of 900, that means he bought it for 3600

OpenStudy (aaronandyson):

Marked price means the price marked on the article during its production

OpenStudy (aaronandyson):

Sale price is the marked price + tax